# Can LLMs Beat Classical Hyperparameter Optimization Algorithms? A Study on autoresearch

### 13 Apr 2026 v4

@@ 2 Related Work
In line with the increasing interest in open-ended agentic discovery (Zhang et al., 2026; Novikov et al., 2025; Lange et al., 2026; Liu et al., 2024; Wang et al., 2026), recent work explores LLMs as components in HPO pipelines. While in principle, methods such as AlphaEvolve (Novikov et al., 2025) or ShinkaEvolve (Lange et al., 2026) can be tasked to optimize the objective in autoresearch, we focus only on methods tailored specifically for HPO. LLAMBO (Ye et al., 2024) uses an LLM as a surrogate model inside Bayesian optimization, replacing the Gaussian process with LLM-based performance predictions. SLLMBO (Mahammadli and Ertekin, 2024) integrates an LLM with TPE into a joint sampler. Zhang et al. (2023) prompt LLMs directly for HP suggestions. LLaMA-ES (Kramer, 2024) uses LLMs to tune CMA-ES’s own hyperparameters. Schwanke et al. (2026a) partition the search space into subregions via a bandit mechanism and use an LLM to propose candidates within each region. The authors further extend that to multi-objective optimization (Schwanke et al., 2026b). However, none of these methods share the classical optimizer’s full internal state with the LLM, which is the key idea behind Centaur.
@@ 3 Experimental Setup
| LLAMBO (Optuna) | Fixed | OptunaHub (Shibata et al., 2025) port: binary surrogate labels, random categorical sampling |
| LLAMBO (Paper) | Fixed | Reimplementation faithful to Ye et al. (2024): continuous labels, all HPs visible |
@@ 5.1 Classical methods outperform LLMs in fixed search spaces
| Method | Trials | Best val_bpb | OOM% | Spread | Pairwise | Step | Cells |
| Centaur | 3345 | 0.97630.0005 | 15% | 0.115 | 0.550 | 0.341 | 330 |
| CMA-ES | 33613 | 0.97850.0036 | 16% | 0.158 | 0.794 | 0.580 | 802 |
| Karpathy Agent (14 HPs) | 2892 | 0.99040.0002 | 1% | 0.035 | 0.225 | 0.057 | 29 |
| Karpathy Agent (Code) | 3247 | 0.98140.0046 | 12% | – | – | – | – |
| LLAMBO (Optuna) | 62964 | 0.98820.0012 | 61% | 0.251 | 1.247 | 1.130 | 458 |
| LLAMBO (Paper) | 4966 | 0.98620.0041 | 48% | 0.252 | 1.257 | 1.218 | 364 |
| Random | 56821 | 0.98730.0021 | 56% | 0.273 | 1.380 | 1.382 | 514 |
| SMAC | 4312 | 0.97780.0020 | 36% | 0.238 | 1.191 | 0.369 | 224 |
| TPE | 31712 | 0.97680.0019 | 11% | 0.196 | 0.996 | 0.413 | 518 |
@@ 5.2 Unconstrained code editing is viable but requires model scale
Scaling the LLM does not consistently close the gap. Comparisons with Gemini 2.5 Flash (Comanici et al., 2025) and Gemini 3.1 Flash-Lite (Google DeepMind, 2026) as the LLM optimizer (Section˜A.3) show no improvement over Qwen3.5-27B. Gemini 3.1 Flash-Lite fails to generate valid code in 87–94% of trials for unconstrained editing, but performs comparably to Qwen3.5-0.8B in the fixed search space, confirming that code editing is where model capability matters most. Gemini 3.1 Pro Preview (Figure˜2) is competitive with Qwen3.5-27B for both unconstrained code editing (0.9826 vs 0.9814) and Centaur (0.9767 vs 0.9763), with lower OOM rates for code editing (2% vs 12%), but does not close the gap to the best classical methods. Claude Opus 4.6 (Anthropic, 2026) narrows the gap for code editing (0.9773 vs 0.9814 with Qwen3.5-27B), making it competitive with CMA-ES (0.9785), but still falls short of TPE (0.9768). For pure LLM code editing, scaling the model helps but does not surpass the best classical methods.
@@ 5.3 Hybrid optimization: best of both worlds
Centaur outperformed all methods including CMA-ES alone by using the LLM on only 30% of trials. As described in Section˜4, the LLM receives CMA-ES’s full internal state (, , ), the top-5 configurations, and the last 20 trials, and almost always overrides CMA-ES’s proposal (100% for 27B, 95% for 0.8B). Despite constituting only 30% of trials, LLM trials contributed 25% of incumbent improvements (Section˜A.6), confirming that the LLM provides complementary value beyond what CMA-ES finds alone. Beyond improving the mean, Centaur substantially reduces CMA-ES’s cross-seed variance: std drops from 0.0036 for CMA-ES alone to 0.0005 for Centaur, suggesting that the LLM stabilizes the optimizer by injecting domain knowledge that prevents unfavorable seeds from drifting. Within the open-weight models tested, Centaur [0.8B] outperformed Centaur [27B], and Gemini 3.1 Pro Preview performed on par with both (0.9766, 0.9763, 0.9767), suggesting a capability plateau where a cheap LLM suffices when paired with a strong classical optimizer. However, Claude Opus 4.6 (Anthropic, 2026) breaks through this plateau: Centaur [Opus 4.6] achieves 0.97390.0012, improving over the open-weight variants by 0.0024 and establishing a new benchmark leader (Figure˜2). This improvement is not explained by better OOM avoidance (17% vs 15% for Qwen 27B) or more trials, but by higher-quality suggestions during the LLM’s 30% of turns. This indicates that the plateau is not fundamental: a sufficiently capable frontier model provides additional gains even in the hybrid setting, though the practical benefit of open-weight models remains that they achieve competitive performance at a fraction of the inference cost. We ablate the LLM ratio (see Appendix Tables˜4 and 11). For the 0.8B model, achieves the best result (0.9735), slightly outperforming the default (0.9776), while degrades to 0.9789. For the 27B model, performs best (0.9745), but collapses to 0.9903, which is worse than CMA-ES alone. This confirms that CMA-ES should retain majority control of the optimization trajectory: the LLM is most effective as an occasional informed perturbation, not as the primary search driver.
@@ A.3 Frontier Model Comparison
Figures˜9 and 10 compare the Flash variants with Qwen3.5-27B, but neither outperforms it. Notably, Karpathy Agent (Code) with Gemini 3.1 Flash-Lite suffers from extremely high failure rates (87–94%), as the model frequently generates code modifications that cause the training script to crash (93% of failures are runtime errors, not OOM). In the fixed search space, however, Gemini 3.1 Flash-Lite performs comparably to Qwen3.5-0.8B. Figure˜2 shows that Gemini 3.1 Pro Preview is competitive with Qwen3.5-27B for both unconstrained code editing (Karpathy Agent Code: 0.98260.0005 vs 0.98140.0057, 2% OOM) and Centaur in the fixed search space (0.97670.0016 vs 0.97630.0006). Gemini Pro achieves lower variance than Qwen for code editing but does not improve the mean, confirming that scaling the LLM optimizer does not close the gap to classical methods.
@@ A.7 Diversity Metrics
The main text focuses on 27B variants; we now report the corresponding 0.8B results. Table˜5 reports diversity metrics for the 0.8B LLM variants of all LLM-based methods.
| Method | Trials | Best val_bpb | OOM% | Spread | Pairwise | Step | Cells |
| Centaur | 3344 | 0.97660.0008 | 15% | 0.133 | 0.651 | 0.390 | 359 |
| Karpathy Agent (14 HPs) | 3383 | 0.99080.0002 | 16% | 0.037 | 0.066 | 0.062 | 24 |
| Karpathy Agent (Code) | 36934 | 0.99100.0001 | 20% | – | – | – | – |
| LLAMBO (Optuna) | 66829 | 0.98730.0003 | 64% | 0.297 | 1.498 | 1.311 | 691 |
| LLAMBO (Paper) | 6056 | 0.98940.0007 | 59% | 0.272 | 1.365 | 1.330 | 290 |
@@ A.11 LLM Prompts and Problem Context
Sources: autoresearch_automl/backends/llambo_original_backend.py and llambo_backend.py. Both variants build their prompts via the LLAMBO surrogate / acquisition templates from Ye et al. [2024], with a custom task description prefixed that we reproduce below. LLAMBO (Optuna) is the [Shibata et al., 2025] port; LLAMBO (Paper) is our faithful reimplementation of the original (details in Section˜A.9).

### 17 Apr 2026 v5

@@ 2 Related Work
In line with the increasing interest in open-ended agentic discovery (Zhang et al., 2026; Novikov et al., 2025; Lange et al., 2026; Liu et al., 2024a; Wang et al., 2026), recent work explores LLMs as components in HPO pipelines. While in principle, methods such as AlphaEvolve (Novikov et al., 2025) or ShinkaEvolve (Lange et al., 2026) can be tasked to optimize the objective in autoresearch, we focus only on methods tailored specifically for HPO. LLAMBO (Liu et al., 2024b) uses an LLM as a surrogate model inside Bayesian optimization, replacing the Gaussian process with LLM-based performance predictions. SLLMBO (Mahammadli and Ertekin, 2024) integrates an LLM with TPE into a joint sampler. Zhang et al. (2023) prompt LLMs directly for HP suggestions. LLaMA-ES (Kramer, 2024) uses LLMs to tune CMA-ES’s own hyperparameters. Schwanke et al. (2026a) partition the search space into subregions via a bandit mechanism and use an LLM to propose candidates within each region. The authors further extend that to multi-objective optimization (Schwanke et al., 2026b). However, none of these methods share the classical optimizer’s full internal state with the LLM, which is the key idea behind Centaur.
@@ 3 Experimental Setup
| LLAMBO (Optuna) | Fixed | OptunaHub (Ozaki et al., 2025) port: binary surrogate labels, random categorical sampling |
| LLAMBO (Paper) | Fixed | Reimplementation faithful to Liu et al. (2024b): continuous labels, all HPs visible |
@@ 5.1 Classical methods outperform LLMs in fixed search spaces
| Method | Trials | Best val_bpb | OOM% | Spread | Step |
| Centaur [Qwen 0.8B] | 3334 | 0.97660.0008 | 13% | 0.131 | 0.368 |
| Centaur [Qwen 27B] | 3345 | 0.97630.0005 | 15% | 0.115 | 0.341 |
| Centaur [Gemini 3.1 Pro] | 34410 | 0.97670.0013 | 20% | 0.135 | 0.496 |
| Centaur [Opus 4.6] | 34013 | 0.97390.0012 | 17% | 0.138 | 0.499 |
| CMA-ES | 33613 | 0.97850.0036 | 16% | 0.158 | 0.580 |
| KA (14 HPs) [Qwen 27B] | 2892 | 0.99040.0002 | 1% | 0.035 | 0.057 |
| KA (Code) [Qwen 0.8B] | 36934 | 0.99100.0001 | 19% | – | – |
| KA (Code) [Qwen 27B] | 3247 | 0.98140.0046 | 12% | – | – |
| KA (Code) [Gemini 3.1 Pro] | 30811 | 0.98260.0004 | 3% | – | – |
| KA (Code) [Opus 4.6] | 3029 | 0.97700.0027 | 5% | – | – |
| LLAMBO (Optuna) [Qwen 27B] | 62964 | 0.98820.0012 | 61% | 0.251 | 1.130 |
| LLAMBO (Paper) [Qwen 27B] | 4966 | 0.98620.0041 | 48% | 0.252 | 1.218 |
| Random | 56821 | 0.98730.0021 | 56% | 0.273 | 1.382 |
| SMAC | 4312 | 0.97780.0020 | 36% | 0.238 | 0.369 |
| TPE | 31712 | 0.97680.0019 | 11% | 0.196 | 0.413 |
@@ 5.2 Unconstrained code editing is viable but requires model scale
Scaling the LLM does not consistently close the gap. Comparisons with Gemini 2.5 Flash (Comanici et al., 2025) and Gemini 3.1 Flash-Lite (Google DeepMind, 2026) as the LLM optimizer (Section˜A.3) show no improvement over Qwen3.5-27B. Gemini 3.1 Flash-Lite fails to generate valid code in 87–94% of trials for unconstrained editing, but performs comparably to Qwen3.5-0.8B in the fixed search space, confirming that code editing is where model capability matters most. Gemini 3.1 Pro Preview (Figure˜2) is competitive with Qwen3.5-27B for both unconstrained code editing (0.9826 vs 0.9814) and Centaur (0.9767 vs 0.9763), with lower OOM rates for code editing (3% vs 12%), but does not close the gap to the best classical methods. Claude Opus 4.6 (Anthropic, 2026) narrows the gap for code editing to 0.97700.0027 (vs 0.9814 with Qwen3.5-27B), making it competitive with CMA-ES (0.9785), but still falls short of TPE (0.97680.0019). Frontier model capability also shows up in the OOM rate (Table˜3): Karpathy Agent (Code)’s failure rate drops sharply with LLM capability (19% for 0.8B, 12% for 27B, 3% for Gemini 3.1 Pro, 5% for Opus 4.6), whereas Centaur’s stays in a narrow range (13%–20%) across model choices. This indicates that for pure code editing, LLM capability directly controls memory-awareness of generated code, while for the hybrid Centaur the CMA-ES side dominates OOM behavior. For pure LLM code editing, scaling the model helps but does not surpass the best classical methods.
@@ 5.3 Hybrid optimization: best of both worlds
Centaur outperformed all methods including CMA-ES alone by using the LLM on only 30% of trials. As described in Section˜4, the LLM receives CMA-ES’s full internal state (, , ), the top-5 configurations, and the last 20 trials, and almost always overrides CMA-ES’s proposal (100% for 27B, 95% for 0.8B). Despite constituting only 30% of trials, LLM trials contributed 25% of incumbent improvements (Section˜A.6), confirming that the LLM provides complementary value beyond what CMA-ES finds alone. Beyond improving the mean, Centaur substantially reduces CMA-ES’s cross-seed variance: std drops from 0.0036 for CMA-ES alone to 0.0005 for Centaur, suggesting that the LLM stabilizes the optimizer by injecting domain knowledge that prevents unfavorable seeds from drifting. Within the open-weight models tested, Centaur [0.8B] outperformed Centaur [27B], and Gemini 3.1 Pro Preview performed on par with both (0.9766, 0.9763, 0.9767), suggesting a capability plateau where a cheap LLM suffices when paired with a strong classical optimizer. However, Claude Opus 4.6 (Anthropic, 2026) breaks through this plateau: Centaur [Opus 4.6] achieves 0.97390.0012, improving over the open-weight variants by 0.0024 and establishing a new benchmark leader (Figure˜2). This improvement is not explained by better OOM avoidance (17% vs 15% for Qwen 27B) or more trials, but by higher-quality suggestions during the LLM’s 30% of turns. This indicates that the plateau is not fundamental: a sufficiently capable frontier model provides additional gains even in the hybrid setting, though the practical benefit of open-weight models remains that they achieve competitive performance at a fraction of the inference cost. We ablate the LLM ratio (see Appendix Tables˜4 and 11). For the 0.8B model, achieves the best result (0.9744), slightly outperforming the default (0.9766), while degrades to 0.9849. For the 27B model, performs best (0.9746), but collapses to 0.9902, which is worse than CMA-ES alone. This confirms that CMA-ES should retain majority control of the optimization trajectory: the LLM is most effective as an occasional informed perturbation, not as the primary search driver.
@@ A.3 Frontier Model Comparison
Figures˜9 and 10 compare the Flash variants with Qwen3.5-27B, but neither outperforms it. Notably, Karpathy Agent (Code) with Gemini 3.1 Flash-Lite suffers from extremely high failure rates (87–94%), as the model frequently generates code modifications that cause the training script to crash (93% of failures are runtime errors, not OOM). In the fixed search space, however, Gemini 3.1 Flash-Lite performs comparably to Qwen3.5-0.8B. Figure˜2 shows that Gemini 3.1 Pro Preview is competitive with Qwen3.5-27B for both unconstrained code editing (Karpathy Agent Code: 0.98260.0004 vs 0.98140.0046, 3% OOM) and Centaur in the fixed search space (0.97670.0013 vs 0.97630.0005). Gemini Pro achieves lower variance than Qwen for code editing but does not improve the mean, confirming that scaling the LLM optimizer does not close the gap to classical methods.
@@ A.7 Diversity Metrics
Table˜5 reports diversity metrics for the remaining 0.8B LLM variants not shown in the main Table˜3, i.e., the fixed-search-space pure LLM methods.
| Method | Trials | Best val_bpb | OOM% | Spread | Step |
| Karpathy Agent (14 HPs) | 3383 | 0.99080.0002 | 16% | 0.037 | 0.062 |
| LLAMBO (Optuna) | 66829 | 0.98730.0003 | 64% | 0.297 | 1.311 |
| LLAMBO (Paper) | 6056 | 0.98940.0007 | 59% | 0.272 | 1.330 |
@@ A.11 LLM Prompts and Problem Context
Sources: autoresearch_automl/backends/llambo_original_backend.py and llambo_backend.py. Both variants build their prompts via the LLAMBO surrogate / acquisition templates from Liu et al. [2024b], with a custom task description prefixed that we reproduce below. LLAMBO (Optuna) is the [Ozaki et al., 2025] port; LLAMBO (Paper) is our faithful reimplementation of the original (details in Section˜A.9).